The decomposition of KClO



3



is commonly used to prepare small amounts of O



2



in the



laboratory:



2 KClO



3



s → 2 KCl s + 3 O



2



g



How many grams of O



2



can be prepared from 4.50 g of KClO



3



?




1
Expert's answer
2021-12-16T13:11:04-0500

The decomposition of KClO3 is commonly used to prepare small amounts of O2 in the laboratory:

2 KClO3(s) → 2 KCl(s) + 3 O2(g)

How many grams of O2 can be prepared from 4.50 g of KClO3?


According to the equation, n (O3) = 3/2 x n (KClO3).

n = m / M

M (O2) = 32 g/mol

M (KClO3) = 122.6 g/mol

n (KClO3) = 4.50 / 122.6 = 0.04 mol

n (O2) = 3/2 x n (KClO3) = 3/2 x 0.04 = 0.06 mol


m (O2) = n x M = 0.06 x 32 = 1.92 g